# Anatomy of an ionized bubble: NIRCam grism spectroscopy of the double-peaked Lyman- emitter COLA1 and its environment

### v1

@@
The increasingly neutral intergalactic gas at impacts the Lyman- (Ly) flux observed from galaxies. One luminous galaxy, COLA1, stands out because of its unique double-peaked Ly line at , unseen in any simulation of reionization. Here we present JWST/NIRCam wide-field slitless spectroscopy in a 21 arcmin2 field centered on COLA1. We find 141 galaxies spectroscopically-selected through the [O III] doublet at , with 40 of these sources showing H. For COLA1 we additionally detect [O III]4363 as well as H. We measure a systemic redshift of for COLA1, confirming the classical double-peak nature of the Ly profile. This implies that it resides in a highly ionized bubble and that it is leaking ionizing photons with a high escape fraction –%, making it a prime laboratory to study Lyman continuum escape in the Epoch of Reionization. COLA1 shows all the signs of a prolific ionizer with a Ly escape fraction of , Balmer decrement indicating no dust, a steep UV slope (), and a star-formation surface density that of typical galaxies at similar redshift. We detect 5 galaxies in COLA1’s close environment (). Exploiting the high spectroscopic completeness inherent to grism surveys, and using mock simulations that fully mimic the selection function, we show the that number of detected companions is very typical for a normal similarly UV-bright () galaxy – that is, the ionized bubble around COLA1 is unlikely due to an excessively large over-density. Instead, the measured ionizing properties suggest that COLA1 by itself might be powering the bubble required to explain its double-peaked Ly profile ( pMpc), with only minor contributions from detected neighbours ().
@@
The Epoch of Reionization (EoR) is the last major phase transition in the Universe. During this period, the bulk of hydrogen gas in the intergalactic medium (IGM) transitioned from neutral to fully ionized. The most recent observations of the Lyman- (Ly; Å) forest in distant quasar spectra indicate that this process was complete at (e.g., Greig & Mesinger, 2017; Bosman et al., 2022), whereas the optical depth to the cosmic microwave background as measured by the Planck satellite places the average reionization redshift around – (Planck Collaboration et al., 2016). The start and duration of the reionization process are still uncertain (see e.g., Robertson, 2022; Gnedin & Madau, 2022), and they are linked to the sources of ionization (e.g., Sharma et al., 2018; Finkelstein et al., 2019; Naidu et al., 2020). Star-formation in galaxies is the leading candidate to have provided the majority of ionizing photons (e.g., Bouwens et al., 2015; Dayal et al., 2020), in particular because UV luminous AGN are very rare at (e.g., Kulkarni et al., 2019), while UV-faint AGN are more common but appear heavily obscured (e.g., Kocevski et al., 2023; Matthee et al., 2023b; Greene et al., 2023; Kokorev et al., 2024).
@@
For star formation, the main uncertainty is whether rare bright or numerous faint galaxies dominated the budget. This hinges upon the luminosity dependence of the escape fraction of ionizing photons (LyC, Å). Detecting LyC photons directly from galaxies becomes increasingly challenging at higher redshifts, as they are attenuated by the increasingly denser and more neutral IGM (Inoue et al., 2014; Steidel et al., 2018). Some semi-empirical and hydrodynamical models suggest that faint () galaxies dominate (e.g., Ocvirk et al., 2021; Trebitsch et al., 2021), others point towards a more important contribution from late-emerging luminous galaxies (in particular supported by the seemingly rapid and late timeline of reionziation; Naidu et al., 2020; Nakane et al., 2023), or towards a more complex picture where the mass of the dominant agent of reionization evolves during the process (Dayal & Giri, 2023) and for most of the time is of intermediate mass (; Rosdahl et al. 2022, see also e.g., Ma et al. 2018).
@@
As the neutral gas in the intergalactic medium absorbs the LyC photons that escaped from galaxies in the epoch of reionization, our knowledge of the amount of LyC leakage and its dependence on the properties of galaxies relies the local and intermediate redshift Universe . The main physical processes that prevent ionizing photon escape are absorption by dust and neutral hydrogen in the interstellar medium (ISM). Likewise, in the – Universe, the dust-sensitive observed UV slope has been recognised as a predictor of LyC escape (e.g., Chisholm et al., 2022), albeit with significant scatter, whereas line ratios that trace the ionization parameter (e.g., [OIII]/[OII]) have yielded inconclusive results (e.g., Izotov et al., 2018; Naidu et al., 2018; Flury et al., 2022). LyC escape fraction further appears to correlate with the star formation rate surface density (e.g., Flury et al., 2022), as this quantity is linked to the presence of large-scale galactic outflows (e.g., Heckman & Borthakur, 2016). Recent studies indicate that high LyC escape is associated with strong UV emission lines as CIV and HeII (e.g., Naidu et al., 2022a; Mainali et al., 2022; Mascia et al., 2023c; Saxena et al., 2022; Schaerer et al., 2022; Kramarenko et al., 2024). This suggests a link between the stellar populations that lead to these high ionisation lines and the processes that enable LyC escape.
@@
The radiative transfer of Ly photons is sensitive to the dust attenuation and the neutral gas column density (e.g., Verhamme et al., 2015; Kakiichi & Gronke, 2021). As these are the same barriers to LyC escape, the strength and shape, in particular the peak separation, of the Ly emission line from galaxies is a powerful indirect tracer of ionizing photon leakage (e.g., Izotov et al., 2018; Steidel et al., 2018; Flury et al., 2022; Pahl et al., 2023). However, the main downside is that measurements of the Ly line-profile as it emerges from the ISM can generally not be obtained for sources in the epoch of reionization. As neutral intergalactic gas in the epoch of reionization impacts the Ly line (e.g., Haiman, 2002; Pentericci et al., 2014; Mason et al., 2018; Gurung-López et al., 2020), we expect to observe lower EW Ly emission, with redder velocity offsets as they reside in increasingly smaller ionized regions beyond (e.g., Hayes & Scarlata, 2023). Indeed, JWST spectroscopy has already revealed such systems with relatively faint and redshifted Ly emission at (e.g., Bunker et al., 2023; Jung et al., 2023). This imprint from the neutral gas in the IGM, which particularly impacts the blue side of the Ly line, tends to hide the information on the ISM properties (and hence, LyC escape) encoded in galaxies’ Ly profiles.
@@
An exception to this are galaxies that reside in such large ionised regions where Ly photons could already have redshifted enough before encountering neutral gas in the IGM such that their emergent profile is visible. Currently a handful of such galaxies are known at (Hu et al., 2016; Songaila et al., 2018; Bosman et al., 2020; Meyer et al., 2021). COLA1, located in the COSMOS field at , is the most luminous galaxy among those with double-peaked Ly line known at . COLA1 has a very narrow peak separation of 220 km s-1 (suggestive of a high LyC escape fraction of %; Verhamme et al. 2017; Izotov et al. 2018) and the only double-peak that has been confirmed by multiple instruments (Keck/DEIMOS, Hu et al. 2016; VLT/X-SHOOTER, Matthee et al. 2018). However, with only the Ly line detected, the systemic redshift was uncertain, allowing alternative explanations for the appearance of the Ly profile, such as a merger (see Matthee et al., 2018, for a discussion).
@@
Besides using these systems to infer their LyC escape fraction from the Ly peak separation, these lines are also suited to probe the sizes of the ionized regions around these galaxies (e.g., Hu et al., 2016; Matthee et al., 2018; Mason & Gronke, 2020). However, sight-lines that allow the detection of double peaks such as the one observed in COLA1 appear extremely rare in hydrodynamical radiative transfer reionization simulations, virtually non existent for luminous galaxies at (Gronke et al., 2021). Therefore, an immediate question that arises is whether such a galaxy resides in a peculiarly large over-density of galaxies, similar to other high-redshift Ly lines (e.g., Witstok et al., 2023).
@@
To measure the systemic redshift and ionizing output of the COLA1 galaxy, together with directly probing galaxies in its environment, we will present results from a survey of COLA1 and its field with JWST/NIRCam imaging and grism Wide Field Slitless Spectroscopy (WFSS) based on a Cycle 1 program (PID 1933, PIs Matthee & Naidu). We observe with the F356W filter, that simultaneously probes the key diagnostic lines H, H and the [O III] doublet for the bright COLA1 galaxy, and (primarily) the strong [O III] doublet in the cMpc environment from . The sensitive imaging in the short-wavelength at 1-2 micron probes the rest-frame UV light of the identified galaxies. The rationale behind our survey design is that the typical strong rest-frame optical emission lines, indicated from Spitzer/IRAC photometry (e.g., Raiter et al., 2010; Labbé et al., 2013; de Barros et al., 2014), would facilitate the direct identification of the galaxies in the environment using the NIRCam grism data. Indeed, spectroscopic observations of large samples of distant galaxies confirm that the strong rest-frame optical lines are common (e.g., Atek et al., 2023; Cameron et al., 2023b; Fujimoto et al., 2023; Matthee et al., 2023a; Oesch et al., 2023; Windhorst et al., 2023). The relatively high resolution of the NIRCam grism, combined with the relatively featureless observed continuum of the major source of light contamination from galaxies at , can indeed be exploited to obtain large samples of spectroscopically selected galaxies with redshifts as accurate as km s-1 (e.g., Kashino et al., 2023).
@@ 2.1.1 Observations
We use imaging and Wide Field Slitless Spectroscopy (WFSS) with NIRCam targeting a mosaic centered around COLA1 through JWST program 1933 (PIs Matthee & Naidu). Observations were performed on January 6 and 7, 2023. The coverage consists of four visits in a mosaic with a position angle of 295 deg, that yields three slightly separated regions of . The total area of the COLA1 field (hereafter C1F) is 21 arcmin2, shown in top panel of Fig. 1.
We use GrismR, which disperses spectra in the horizontal direction of the detectors, in opposite directions for modules A and B. This is similar to the strategy employed by the EIGER survey (Kashino et al., 2023), albeit here maximising the area with cross-dispersed coverage at the expense of gaps. The central are covered by four visits. Another area of 11.4 arcmin2 is covered by two visits and one of 6.2 arcmin2 by a single visit. Grism spectroscopy inherently obtains a 2D spectrum for every source in the field of view, without target pre-selection. However, due to this extensive coverage, the spectra are prone to be contaminated due to blending of spectra from different sources, and the background is also dispersed, adding noise.
@@
Simultaneously with the deep grism exposures, we obtain imaging data in the F115W and F200W filters in the same part of sky. Direct and out of field images are taken using the F150W and F356W filters to facilitate the identification of all spectral traces on the grism data. The total spectroscopic on sky integration time ranges from 8.5 to 34 ks for the F356W/GrismR data, depending on the position in the observed field and according to the number of visits as described above. The F115W and F200W imaging was equal to 42 and 58 % of the F356W/GrismR exposure time, respectively. Direct and out of field images amount to 1.5–6 ks of exposure time in the F150W and F356W imaging.
@@
2.1.2 Data reduction & Quality
@@
For the imaging data, we first use the standard Detector1 and Image2 steps and then align the images to the astrometry from stars and spherically symmetric objects (i.e. whose centroid is unambiguously defined in our high resolution images) in the COSMOS2020 catalog (Weaver et al., 2022), which has been aligned to GAIA. We use this indirect method as only few GAIA-detected stars are present in our NIRCam images. By creating source-masked pixel-based median stacks of the imaging in a specific filter/module/camera combination, we subtract the sky and stray-light features (so-called wisps). Cosmic-ray related artefacts (snowballs) are masked following Merlin et al. (2022). The read noise in our imaging data is subtracted using the median sky in quarter rows, columns and in the four amplifiers, respectively. Finally, images are combined with Image3 to a common grid with pixel scale of 0.03″.
@@ 3 Methods
In this section, first we describe the methods used to identify the [O III] emitter sample in the field, and measure line fluxes and systemic redshifts from the grism data. Secondly, we describe the SED modeling code used to infer properties of the sample from the photometric and spectroscopic data. Finally we introduce mock catalogs of the field, employed to characterize the environment of COLA1.
@@
The WFSS technique yields a spectrum for all the objects in the spectroscopic field of view that is shown in Fig. 1. We extract 2D spectra from the GrismR (SCI, EMLINE and CONT) images for the sample of all F356W detected sources (following the so-called ‘forward’ method outlined in Kashino et al. 2023). This sample typically contains sources down to a F356W magnitude 28 but extends to magnitude 29.5 in the deepest areas. We run SExtractor (Bertin & Arnouts, 1996) on each coadded EMLINE 2D spectrum to identify pairs of emission lines, regardless of any photometric pre-selection apart from detection. Then, we select an initial [O III] doublet candidate sample by identifying pairs of lines with S/N , with wavelengths that could be compatible with [O III]λ4960,λ5008 at –, with a maximum offset of 3 pixels (0.09″) in the spatial centroid (i.e., perpendicular to the spectral direction). We require a peak separation for the [O III] doublet according to the initial redshift solution, within a 15 Å tolerance. The theoretical line flux ratio of the [O III] doublet is (Storey & Zeippen, 2000), therefore we conservatively impose a line flux ratio of . This search produces an initial sample of 3221 candidates.
@@
Some of our candidates are multiple-component systems that are sometimes blended into a single catalog entry, sometimes not (see Figures 15 and 16 for some examples). Following Matthee et al. (2023a), we define a ‘system’ as a group of two or more galaxies with a projected separation smaller than 2″ and an absolute redshift difference of km s-1. In such systems, the measured line flux of each single component is possibly contaminated by their companions. Moreover, whether the components of systems can disentangled or not is an effect of our arbitrary line of sight to the system a the spatial resolution of the images. We find three such systems within our list of galaxies. We define the main component of each system as the brightest galaxy in F356W, and add the photometry and line fluxes of their companion to the main component in the catalog. The companions are then removed from our galaxy catalog, and the systems are regarded as single objects, leaving a sample of 141 [O III] emitters. Among this sample, 40 objects () have a measurement of H flux with S/N (13 with S/N ).
A complete list of all the selected sources is found in Table 4, and some examples of their grism spectra are shown in Appendix C. In order to carry out this task, we created inspection images similar to Fig. 2. Two of the authors (ATT, JM) agreed on the final sample. In order to visually select our candidates, both peaks of the [O III] doublet must be present in the available grism modules. The shape of these emission lines must not substantially differ from the morphology seen in the detection image, nor between modules. It has to be taken into account that, since modules A and B disperse photons in opposite directions, the transverse spatial component is mirrored between modules in our extraction method. Also, both peaks of the doublet must not be significantly misaligned in the spatial direction, nor with the detection image. Additionally, color is another useful indicator for identifying contaminants. In the color representation we chose, [O III] emitters present red-purple colors, as a result of typically blue colors in (i.e. their UV continuum), and [O III]-boosted red colors in . Sometimes an [O III] doublet is assigned to more than one object, due to the intrinsic degeneracy of the grism wavelength solution – galaxies that line up along the dispersion direction end up with emission lines in the same row. In those cases we assign the correct object matching the morphologies of the grism emission lines and photometric image.

### v2

@@
The increasingly neutral intergalactic gas at impacts the Lyman- (Ly) flux observed from galaxies. One luminous galaxy, COLA1, stands out because of its unique double-peaked Ly line at , unseen in any simulation of reionization. Here, we present JWST/NIRCam wide-field slitless spectroscopy in a 21 arcmin2 field centered on COLA1. We find 141 galaxies spectroscopically selected through the [O III] doublet at , with 40 of these sources showing H. For COLA1, we additionally detect [O III]4363 as well as H. We measure a systemic redshift of for COLA1, confirming the classical double-peak nature of the Ly profile. This implies that it resides in a highly ionized bubble and that it is leaking ionizing photons with a high escape fraction of –%, making it a prime laboratory to study Lyman continuum escape in the Epoch of Reionization. COLA1 shows all the signs of a prolific ionizer with a Ly escape fraction of , Balmer decrement indicating no dust, a steep UV slope (), and a star-formation surface density that of typical galaxies at similar redshift. We detect five galaxies in COLA1’s close environment (). Exploiting the high spectroscopic completeness inherent to grism surveys, and using mock simulations that fully mimic the selection function, we show that the number of detected companions is very typical for a normal similarly UV-bright () galaxy – that is, the ionized bubble around COLA1 is unlikely to be due to an excessively large over-density. Instead, the measured ionizing properties suggest that COLA1 by itself might be powering the bubble required to explain its double-peaked Ly profile ( pMpc), with only minor contributions from detected neighbors ().
@@
The Epoch of Reionization (EoR) is the last major phase transition in the Universe. During this period, the bulk of hydrogen gas in the intergalactic medium (IGM) transitioned from neutral to fully ionized. The most recent observations of the Lyman- (Ly; Å) forest in distant quasar spectra indicate that this process was complete at (e.g., Greig & Mesinger, 2017; Bosman et al., 2022), whereas the optical depth to the cosmic microwave background as measured by the Planck satellite places the average reionization redshift around – (Planck Collaboration et al., 2016). The start and duration of the reionization process are still uncertain (see e.g., Robertson, 2022; Gnedin & Madau, 2022), and they are linked to the sources of ionization (e.g., Sharma et al., 2018; Finkelstein et al., 2019; Naidu et al., 2020). Star formation in galaxies is the leading candidate to have provided the majority of ionizing photons (e.g., Bouwens et al., 2015; Dayal et al., 2020), in particular because ultraviolet (UV) luminous active galactic nuclei (AGN) are very rare at (e.g., Kulkarni et al., 2019), while UV-faint AGN are more common but appear heavily obscured (e.g., Kocevski et al., 2023; Matthee et al., 2024; Greene et al., 2024; Kokorev et al., 2024).
@@
For star formation, the main uncertainty is whether rare bright or numerous faint galaxies dominated the budget. This hinges upon the luminosity dependence of the escape fraction of ionizing photons (LyC, Å). Detecting LyC photons directly from galaxies becomes increasingly challenging at higher redshifts, as they are attenuated by the increasingly denser and more neutral IGM (Inoue et al., 2014; Steidel et al., 2018). Some semiempirical and hydrodynamical models suggest that faint () galaxies dominate (e.g., Ocvirk et al., 2021; Trebitsch et al., 2021); others point toward a more important contribution from late-emerging luminous galaxies (in particular supported by the seemingly rapid and late timeline of reionziation; Naidu et al., 2020; Nakane et al., 2024), or toward a more complex picture in which the mass of the dominant agent of reionization evolves during the process (Dayal & Giri, 2024) and for most of the time is of intermediate mass (; Rosdahl et al. 2022, see also e.g., Ma et al. 2018).
@@
As the neutral gas in the IGM absorbs the LyC photons that escaped from galaxies in the EoR, our knowledge of the amount of LyC leakage and its dependence on the properties of galaxies relies on the local and intermediate redshift Universe at . The main physical processes that prevent ionizing photon escape are absorption by dust and neutral hydrogen in the interstellar medium (ISM). Likewise, in the – Universe, the dust-sensitive observed UV slope has been recognized as a predictor of LyC escape (e.g., Chisholm et al., 2022), albeit with significant scatter, whereas line ratios that trace the ionization parameter (e.g., [OIII]/[OII]) have yielded inconclusive results (e.g., Izotov et al., 2018; Naidu et al., 2018; Flury et al., 2022). The LyC escape fraction further appears to correlate with the star formation rate surface density (e.g., Flury et al., 2022), as this quantity is linked to the presence of large-scale galactic outflows (e.g., Heckman & Borthakur, 2016). Recent studies indicate that high LyC escape is associated with strong UV emission lines as CIV and HeII (e.g., Naidu et al., 2022a; Mainali et al., 2022; Mascia et al., 2023b; Saxena et al., 2022; Schaerer et al., 2022; Kramarenko et al., 2024). This suggests a link between the stellar populations that lead to these high ionization lines and the processes that enable LyC escape.
@@
The radiative transfer of Ly photons is sensitive to the dust attenuation and the neutral gas column density (e.g., Verhamme et al., 2015; Kakiichi & Gronke, 2021). As these are the same barriers to LyC escape, the strength and shape, in particular the peak separation, of the Ly emission line from galaxies is a powerful indirect tracer of ionizing photon leakage (e.g., Izotov et al., 2018; Steidel et al., 2018; Flury et al., 2022; Pahl et al., 2023). However, the main downside is that measurements of the Ly line-profile as it emerges from the ISM can generally not be obtained for sources in the EoR. As neutral intergalactic gas in the EoR impacts the Ly line (e.g., Haiman, 2002; Pentericci et al., 2014; Mason et al., 2018; Gurung-López et al., 2020), we expect to observe lower EW Ly emission, with redder velocity offsets as they reside in increasingly smaller ionized regions beyond (e.g., Hayes & Scarlata, 2023). Indeed, JWST spectroscopy has already revealed such systems with relatively faint and redshifted Ly emission at (e.g., Bunker et al., 2023; Jung et al., 2024). This imprint from the neutral gas in the IGM, which particularly impacts the blue side of the Ly line, tends to hide the information on the ISM properties (and hence, LyC escape) encoded in galaxies’ Ly profiles.
@@
An exception to this are galaxies that reside in such large ionized regions where Ly photons could already have redshifted enough before encountering neutral gas in the IGM such that their emergent profile is visible. Currently, a handful of such galaxies are known at (Hu et al., 2016; Songaila et al., 2018; Bosman et al., 2020; Meyer et al., 2021). COLA1, located in the COSMOS field at , is the most luminous galaxy among those with double-peaked Ly line known at . COLA1 has a very narrow peak separation of 220 km s-1 (suggestive of a high LyC escape fraction of %; Verhamme et al. 2017; Izotov et al. 2018) and the only double-peak that has been confirmed by multiple instruments (Keck/DEIMOS, Hu et al. 2016; VLT/X-shooter, Matthee et al. 2018). However, with only the Ly line detected, the systemic redshift was uncertain, allowing alternative explanations for the particular shape of the Ly profile, such as a merger (see Matthee et al., 2018, for a discussion).
@@
Besides using these systems to infer their LyC escape fraction from the Ly peak separation, these lines are also suited to probe the sizes of the ionized regions around these galaxies (e.g., Hu et al., 2016; Matthee et al., 2018; Mason & Gronke, 2020). However, sight-lines that allow the detection of double peaks such as the one observed in COLA1 appear extremely rare in hydrodynamical radiative transfer reionization simulations, virtually non existent for luminous galaxies at (Gronke et al., 2021). Therefore, an immediate question that arises is whether such a galaxy resides in a peculiarly large over-density of galaxies, similar to other high-redshift Ly lines (e.g., Witstok et al., 2024).
@@
To measure the systemic redshift and ionizing output of the COLA1 galaxy, together with directly probing galaxies in its environment, we will present results from a survey of COLA1 and its field with JWST/NIRCam imaging and grism Wide Field Slitless Spectroscopy (WFSS) based on a Cycle 1 program (PID 1933, PIs Matthee & Naidu). We observe with the F356W filter, which simultaneously probes the key diagnostic lines H, H and the [O III] doublet for the bright COLA1 galaxy, and (primarily) the strong [O III] doublet in the cMpc environment from . The sensitive imaging in the short-wavelength at 1-2 micron probes the rest-frame UV light of the identified galaxies. The rationale behind our survey design is that the typical strong rest-frame optical emission lines, indicated from Spitzer/IRAC photometry (e.g., Raiter et al., 2010; Labbé et al., 2013; de Barros et al., 2014), would facilitate the direct identification of the galaxies in the environment using the NIRCam grism data. Indeed, spectroscopic observations of large samples of distant galaxies confirm that the strong rest-frame optical lines are common (e.g., Atek et al., 2023; Cameron et al., 2023b; Fujimoto et al., 2023; Matthee et al., 2023; Oesch et al., 2023; Windhorst et al., 2023). The relatively high resolution of the NIRCam grism, combined with the relatively featureless observed continuum of the major source of light contamination from galaxies at , can indeed be exploited to obtain large samples of spectroscopically selected galaxies with redshifts as accurate as km s-1 (e.g., Kashino et al., 2023).
@@ 2.1.1 Observations
We use imaging and Wide Field Slitless Spectroscopy (WFSS) with NIRCam targeting a mosaic centered around COLA1 through JWST program 1933 (PIs Matthee & Naidu). Observations were performed on January 6 and 7, 2023. The coverage consists of four visits in a mosaic with a position angle of 295 deg, which yields three slightly separated regions of . The total area of the COLA1 field (hereafter C1F) is 21 arcmin2, shown in top panel of Fig. 1.
We use GrismR, which disperses spectra in the horizontal direction of the detectors, in opposite directions for modules A and B. This is similar to the strategy employed by the EIGER survey (Kashino et al., 2023), albeit here maximizing the area with cross-dispersed coverage at the expense of gaps. The central are covered by four visits. Another area of 11.4 arcmin2 is covered by two visits and one of 6.2 arcmin2 by a single visit. Grism spectroscopy inherently obtains a 2D spectrum for every source in the field of view, without target preselection. However, due to this extensive coverage, the spectra are prone to be contaminated due to blending of spectra from different sources, and the background is also dispersed, adding noise.
@@
Simultaneously with the deep grism exposures, we obtain imaging data in the F115W and F200W filters in the same part of the sky. Direct and out of field images are taken using the F150W and F356W filters to facilitate the identification of all spectral traces on the grism data. The total spectroscopic on sky integration time ranges from 8.5 to 34 ks for the F356W/GrismR data, depending on the position in the observed field and according to the number of visits as described above. The F115W and F200W imaging was equal to 42 and 58 % of the F356W/GrismR exposure time, respectively. Direct and out of field images amount to 1.5–6 ks of exposure time in the F150W and F356W imaging.
@@
2.1.2 Data reduction and quality
@@
For the imaging data, we first use the standard Detector1 and Image2 steps and then align the images to the astrometry from stars and spherically symmetric objects (i.e., whose centroid is unambiguously defined in our high resolution images) in the COSMOS2020 catalog (Weaver et al., 2022), which has been aligned to GAIA. We use this indirect method as only few GAIA-detected stars are present in our NIRCam images. By creating source-masked pixel-based median stacks of the imaging in a specific filter/module/camera combination, we subtract the sky and stray-light features (so-called wisps). Cosmic-ray related artefacts (snowballs) are masked following Merlin et al. (2022). The read noise in our imaging data is subtracted using the median sky in quarter rows, columns and in the four amplifiers, respectively. Finally, images are combined with Image3 to a common grid with pixel scale of 0.03″.
@@ 3 Methods
In this section, first we describe the methods used to identify the [O III] emitter sample in the field (Sect. 3.1), and measure line fluxes and systemic redshifts from the grism data (Sects. 3.2 and 3.3). Secondly, we describe the SED modeling code used to infer properties of the sample from the photometric and spectroscopic data (Sect. 3.4). Finally we introduce mock catalogs of the field, employed to characterize the environment of COLA1 (Sect. 3.5).
@@
The WFSS technique yields a spectrum for all the objects in the spectroscopic field of view that is shown in Fig. 1. We extracted 2D spectra from the GrismR (SCI, EMLINE and CONT) images for the sample of all F356W detected sources (following the so-called “forward” method outlined in Kashino et al. 2023). This sample typically contains sources down to a F356W magnitude of 28 but extends to magnitude of 29.5 in the deepest areas. We ran SExtractor (Bertin & Arnouts, 1996) on each coadded EMLINE 2D spectrum to identify pairs of emission lines, regardless of any photometric preselection apart from detection. Then, we selected an initial [O III] doublet candidate sample by identifying pairs of lines with S/N , with wavelengths that could be compatible with [O III]λ4960,λ5008 at –, with a maximum offset of 3 pixels (0.09″) in the spatial centroid (i.e., perpendicular to the spectral direction). We required a peak separation for the [O III] doublet according to the initial redshift solution, within a 15 Å tolerance. The theoretical line flux ratio of the [O III] doublet is (Storey & Zeippen, 2000), and therefore we conservatively imposed a line flux ratio of . This search produced an initial sample of 3221 candidates.
@@
Some of our candidates are multiple-component systems that are sometimes blended into a single catalog entry, sometimes not (see Fig. 15 for some examples). Following Matthee et al. (2023), we define a ‘system’ as a group of two or more galaxies with a projected separation smaller than 2″ and an absolute redshift difference of km s-1. In such systems, the measured line flux of each single component is possibly contaminated by their companions. Moreover, whether the components of systems can disentangled or not is an effect of our arbitrary line of sight to the system a the spatial resolution of the images. We find three such systems within our list of galaxies. We define the main component of each system as the brightest galaxy in F356W, and add the photometry and line fluxes of their companion to the main component in the catalog. The companions are then removed from our galaxy catalog, and the systems are regarded as single objects, leaving a sample of 141 [O III] emitters. Among this sample, 40 objects () have a measurement of H flux with S/N (13 with S/N ).
A complete list of all the selected sources is found in Table 4, and some examples of their grism spectra are shown in Appendix C. In order to carry out this task, we created inspection images similar to Fig. 2. Two of the authors (ATT, JM) agreed on the final sample. In order to visually select our candidates, both peaks of the [O III] doublet must be present in the available grism modules. The shape of these emission lines must not substantially differ from the morphology seen in the detection image, nor between modules. It has to be taken into account that, since modules A and B disperse photons in opposite directions, the transverse spatial component is mirrored between modules in our extraction method. Also, both peaks of the doublet must not be significantly misaligned in the spatial direction, nor with the detection image. Additionally, color is another useful indicator for identifying contaminants. In the color representation we chose, [O III] emitters present red-purple colors, as a result of typically blue colors in (i.e., their UV continuum), and [O III]-boosted red colors in . Sometimes an [O III] doublet is assigned to more than one object, due to the intrinsic degeneracy of the grism wavelength solution – galaxies that line up along the dispersion direction end up with emission lines in the same row. In those cases we assign the correct object matching the morphologies of the grism emission lines and photometric image.